two bodies are thrown from same point with the same velocity of 50m/s . if their angles of projection are complementary to each other and the difference of maximum heights is 30m,the minimum and maximum heights are?
  1. 50m and 80m
  2. 47.5m and 77.5 m
  3. 30m and 60m
  4. 25m and 55 m

Grade:Select Grade

1 Answers

Priya
6 years ago
let the angle made with horizontal by 1st body is theta.
so,  the angle made with horizontal by 2nd body is (90 - theta).
h1 + h2 = u2sin2theta/2g +  u2sin2(90-theta)/2g
            = u2sin2theta/2g +  u2cos2theta/2g
            = u2/2g (sin2theta+cos2theta)
            = u2/2g
            = 50 * 50 / 2*10
            = 125......(1)
h1-h2    = 30........(2)
subtacting (2) from (1),
h2 = 47.5 and h1= 77.5